OpenStudy (anonymous):

Suppose f(x) = x^2 - 1 and g(x) = 2x. Find f(g(3)).

OpenStudy (anonymous):

If g(x) = 2x then to find f(g(x)) just plug 2x in for x in your f(x) function.

OpenStudy (anonymous):

Just like you'd plug a in for x to find f(a)

OpenStudy (anonymous):

so what would the answer be?

OpenStudy (anonymous):

35

OpenStudy (anonymous):

\[g(3) = 2(3) = 6\]\[\implies f(g(3)) = f(6) = 6^2 -1 = 35\]

OpenStudy (anonymous):

^^ is that correct?

OpenStudy (anonymous):

thank you